# GranQ: Granular Zero-Shot Quantization with Channel-Wise Activation Scaling in QAT

### v4

@@ Abstract
Zero-shot quantization (ZSQ) enables neural network compression without original training data, making it a promising solution for restricted data access scenarios. To compensate for the lack of data, recent ZSQ methods typically rely on synthetic inputs generated from the full-precision model. However, these synthetic inputs often lead to activation distortion, especially under low-bit settings. As a result, existing methods struggle to mitigate this issue due to coarse activation scaling. To address this issue, we propose GranQ, a novel activation quantization framework that efficiently applies per-channel scaling through vectorized computation. In contrast to conventional channel-wise methods, which apply vectorization only to the quantization step, GranQ improves efficiency by vectorizing the scaling operation. This design allows GranQ to maintain fine-grained quantization granularity with minimal computational overhead, even in low-bit environments. Extensive experiments under quantization-aware training (QAT) settings demonstrate that GranQ consistently outperforms state-of-the-art ZSQ methods across CIFAR and ImageNet. In particular, our method achieves up to 5.45% higher accuracy in the 3-bit setting on CIFAR-100 and even surpasses the full-precision baseline on CIFAR-10. Furthermore, GranQ achieves significant speedup in quantization latency over conventional per-channel methods, demonstrating improved efficiency. With these findings, we anticipate that GranQ will inspire future research beyond conventional ZSQ approaches centered on data generation and model fine-tuning.
1 Introduction
@@
Neural network compression has been extensively studied for the practical deployment of large-scale deep learning (DL) models. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression include quantization [1, 2, 3], pruning [4, 5, 6, 7], knowledge distillation [8, 9, 10], and neural architecture search [11, 12], as surveyed in [13, 14]. Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model. However, it often requires fine-tuning or calibration to match full-precision (FP) model performance [3]. To address this, zero-shot quantization (ZSQ), also known as data-free quantization, has been proposed to compress models without the original training data. Since the introduction of ZeroQ [15], studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
First, data generation studies focus on generating high-quality data to effectively train the Q model [16, 17, 18]. Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training [19, 20]. Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation [21, 22, 23, 24].
Based on this analysis, we introduce GranQ, a novel ZSQ approach that enables channel-wise activation quantization by adapting the scale to each channel of every input activation. This allows us to overcome the limitations of conventional activation quantization. This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, We apply vectorization to the scaling operation, which is not applied in conventional channel-wise quantization, reducing latency while enabling fine-grained activation quantization. Our contributions can be summarized as follows:
(a) Layer-wise Quantization
(b) GranQ (Channel-wise, Ours)
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. We introduce a fine-grained activation quantization mechanism that considers every channel to reduce activation loss. Although per-channel activation quantization is known to improve precision, it has not been widely adopted due to its high computational cost. We address this by applying vectorization to the scaling operation, enabling efficient per-channel quantization with minimal overhead. To the best of our knowledge, our approach is the first to address the ZSQ problem.
2 Related Work
2.1 Quantization
Quantization reduces the representational range of deep neural networks (DNNs) and minimizes memory usage. It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) [3, 25]. PTQ applies quantization after training without further updates, which makes it efficient and easy to implement. However, it is sensitive to scaling errors and often relies on calibration to improve accuracy. In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator [3, 26]. Both methods require data—PTQ for calibration and QAT for fine-tuning. In practice, access to training data is often limited. To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data [16].
@@ 2.1 Quantization
ZeroQ (CVPR 20) [15]
GDFQ (ECCV 20) [16]
DSG (CVPR 21) [27]
MixMix (CVPR 21) [28]
Qimera (NeurIPS 21) [29]
IntraQ (CVPR 22) [30]
GENIE (CVPR 23) [31]
AdaDFQ (CVPR 23) [17]
Casual-DFQ (ICCV 23) [32]
TexQ (NeurIPS 23) [33]
RIS (AAAI 24) [34]
GenQ (ECCV 24) [18]
SQuant (ICLR 22) [19]
UDFC (ICCV 23) [20]
AIT (CVPR 22) [21]
PLF (CVPR 24) [22]
SynQ (ICLR 25) [24]
@@
Table 1: Categorizing zero-shot quantization algorithms. PTQ denotes post-training quantization, and QAT denotes quantization-aware training. Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
2.2 Zero-shot Quantization
@@
As summarized in Table 1, ZSQ was initially introduced by ZeroQ [15], leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data [31, 18] and developing more effective methods to utilize them [21, 23, 24]. However, current ZSQ methods heavily rely on augmented synthetic inputs. This causes large variation in activation scales across channels. In practice, such input-dependent channel-wise quantization is rarely adopted due to its high computational cost. Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings [35], or exclusively in PTQ scenarios with access to a small amount of data [36, 37]. This highlights the need for efficient channel-wise activation quantization for QAT in ZSQ.
3 Preliminaries and Problem Definition
@@
3.1 Activation Quantization
Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers [38]. It is commonly used with weight quantization to compress models, and the linear quantization scheme is most widely adopted. This process typically consists of quantization, scaling parameter computation, and dequantization [1, 3, 39].
@@ 3.1 Activation Quantization
where $x_{\min}$ and $x_{\max}$ denote the range of activation values, and $b$ is the bit-width. A large range yields a large $s$, increasing quantization error; a smaller range allows finer resolution.
During inference, the dequantization step restores quantized values to real numbers:
$x_{\text{deq}}=s(x_{q}-z)$ (3)
This enables recovery of activation information needed for final predictions.
3.2 Problem Definition
@@
(Problem 2) Severe Quantization Error in Certain Layers
During activation quantization, certain layers induce severe quantization errors, leading to significant activation distortions. These distortions substantially differ from the original distribution and can directly impact the performance of the model.
4 Observation and Methodology
We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce the proposed method known as GranQ.
Figure 2: Activation magnitudes of ResNet-20 on CIFAR-10 using 32-bit FP (left) and 3-bit layer-wise quantization (right). Severe distortions appear, especially in the final layer.
4.1 Observation
Layer-wise quantization effectively captures the activation distribution of each layer. However, it becomes inefficient when activation distributions within a layer exhibit significant variation along the channel axis. Particularly in ZSQ, as shown in the FP graph in Figure 1(a), the activation distributions can differ substantially across channels even within the same layer. This implies that limited representation bits are forced to capture a wide range of activation values, making fine-grained quantization difficult. It consistently appears across all layers. We analyzed this phenomenon occurring across all layers and provided a summary in Table 2. It presents the cosine similarity $\frac{X\cdot Q}{|X|_{2}|Q|_{2}}$, along with the relative error $\frac{|X-Q|_{2}}{|X|_{2}}$. From the results, we observe that layer-wise quantization fails to preserve activation information effectively. This reveals that using a single scaling range for activation quantization has a critical limitation, particularly in ZSQ. In contrast, GranQ, which applies scaling per channel, achieves 1.34$\times$ higher similarity and 2.94$\times$ lower relative error than layer-wise quantization, demonstrating its effectiveness in preserving activation information.
Furthermore, Figure 2 provides a visual representation of the effect of quantization on activation values across layers. In layer-wise quantizaiton, although the overall activations generally reflect the FP activations to some extent, certain layers exhibit fine activation values being mapped to discrete values, leading to coarse representations. These coarse representations have been identified as a major factor leading to a decline in the overall expressiveness and inference performance of the model. Further analysis of Figure 2 is provided in Figure 5 and in Appendix (Figure 7).
@@ 4.1 Observation
Table 2: Average cosine similarity and relative error for 3-bit activation quantization on ResNet-20 (CIFAR-100). GranQ (channel-wise) outperforms layer-wise quantization by achieving less distortion in activation representation, as indicated by higher similarity and lower error.
4.2 Methodology
Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead (Figure 3). GranQ applies channel-wise scaling through activation reshaping and decomposition, and achieves fast parallel processing via vectorization of scaling parameters, which enables efficient computation of both scaling and quantization.
4.2.1 Activation Reshaping and Decomposition
We propose activation reshaping and decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$. This transformation enables decomposition of the activation map along the channel axis, allowing each channel to be processed independently. It also facilitates vectorized computation of per-channel statistics (e.g., min and max) required for scaling factor calculation during quantization. The activation reshaping and decomposition is defined as follows:
@@ 4.2.1 Activation Reshaping and Decomposition
4.2.2 Vectorization of Scaling Parameters
In the vectorization of scaling parameters stage, multiple scaling ranges are stored in a vectorized form for each channel. This allows dynamic adaptation to the diverse activation distributions induced by synthetic inputs. Thus, GranQ contrasts with traditional per-channel quantization, which uses fixed scaling parameters for each channel regardless of input variation. As shown in Equation 5, the scaling factors $\vec{s}$ and zero-points $\vec{z}$ are computed in vectorized form. The activation values are then quantized and dequantized according to Equation 6, with all operations applied in parallel across channels.
@@ 4.2.2 Vectorization of Scaling Parameters
$A_{q,l}=\left\lfloor\frac{A}{\vec{s}}+\vec{z}\right\rceil,A_{\text{deq},l}=\vec{s}(A_{q,l}-\vec{z})$ (6)
This process allows efficient and adaptive quantization of activation tensors exhibiting high variation across inputs, leading to improved computational efficiency (Table 7 in appendix).
[…]
5 Experiments
@@
5.1 Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 [40], and ImageNet (ILSVRC 2012) [41] validation datasets. For the CIFAR datasets, ResNet-20 [42] was used as the quantization model, whereas ResNet-18 [42], ResNet-50 [42], and MobileNetV2 [43] were employed for ImageNet. All experiments were conducted using the SGD optimizer [44] with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods [16, 21, 17, 33, 34, 22, 18, 23, 24]. For data generation, we followed the AdaDFQ [17] approach based on ACGAN [45]. Further implementation details are provided in the appendix, Section A.3.2 and A.3.3. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
@@ 5.1 Experimental Setup and Details
[…]
@@ 5.2 Performance Evaluation
CIFAR-10/100. As indicated in Table 3, GranQ consistently achieved the highest accuracy across all bitwidths. For CIFAR-10, it attained accuracies of 94.06% (5-bit), 93.52% (4-bit), and 91.37% (3-bit). For CIFAR-100, the results were 70.05% (5-bit), 68.79% (4-bit), and 62.73% (3-bit). Notably, GranQ outperformed SynQ [24], the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting. This result demonstrates its ability to effectively overcome the limitations of conventional low-bitwidth quantization techniques. Overall, GranQ consistently outperformed existing methods across all bitwidths in both CIFAR-10 and CIFAR-100, with particularly strong improvements in the 3-bit quantization setting. Remarkably, in the CIFAR-10 5-bit setting, GranQ even surpassed the performance of the FP model. These results suggest that we can effectively apply GranQ to small- and medium-scale datasets.
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model [46] and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods [16]. In contrast, GranQ improves the quantization mechanism without relying heavily on diffusion-based data generation, as GenQ does. As our method is adaptable to existing QAT frameworks, it can potentially achieve even higher performance when combined with GenQ’s data generation.
5.3 Ablation Study
5.3.1 Effectiveness Evaluation
@@
First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ [16] and AdaDFQ [17]. The integration of GranQ into GDFQ [16] led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting. Similarly, AdaDFQ [17] exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%). These results suggest that GranQ effectively reduces quantization errors when combined with data synthesis-based quantization methods, leading to enhanced model performance.
Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT [21] and AdaDFQ+AKT [23]. For Qimera+AIT [21], the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%). Similarly, for AdaDFQ+AKT [23], the accuracy improved from 54.68% to 62.01% (+7.33%). These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
5.3.2 Efficiency Evaluation
@@
Layer-wise 49.98 ($\pm$ 0.24) 0.227
Channel-wise (Conventional scalar-loop) 62.68 ($\pm$ 0.11) 103.671
Channel-wise (Vectorized, GranQ) 62.73 ($\pm$ 0.17) 0.236

### v5

@@ Abstract
Zero-shot quantization (ZSQ) enables neural network compression without original training data, making it a promising solution for restricted data access scenarios. To compensate for the lack of data, recent ZSQ methods typically rely on synthetic inputs generated from the full-precision model. However, these synthetic inputs often lead to activation distortion, especially under low-bit settings. To mitigate this, existing methods typically employ per-channel scaling, but they still struggle due to the severe computational overhead during the accumulation process. To overcome this critical bottleneck, we propose GranQ, a novel activation quantization framework that introduces an efficient pre-scaling strategy. Unlike conventional channel-wise methods that repeatedly perform scaling operations during accumulation, GranQ applies scaling factors in a pre-scaling step through fully vectorized computation, eliminating runtime scaling overhead. This design enables GranQ to maintain fine-grained quantization accuracy while significantly reducing computational burden, particularly in low-bit quantization settings. Extensive experiments under quantization-aware training (QAT) settings demonstrate that GranQ consistently outperforms state-of-the-art ZSQ methods across CIFAR and ImageNet. In particular, our method achieves up to 5.45% higher accuracy in the 3-bit setting on CIFAR-100 and even surpasses the full-precision baseline on CIFAR-10. Furthermore, GranQ achieves significant speedup in quantization latency over conventional per-channel methods, demonstrating improved efficiency. With these findings, we anticipate that GranQ will inspire future research beyond conventional ZSQ approaches centered on data generation and model fine-tuning. The official code is available at https://github.com/anonymus-orange/GranQ.
Introduction
@@
Neural network compression has been extensively studied for the practical deployment of large-scale deep learning (DL) models. In particular, reducing model size while minimizing performance degradation is crucial for utilizing DL on edge devices (e.g., mobile phones, embedded systems, and drones). Major approaches to model compression include quantization (Hubara et al. 2018; Li et al. 2021; Gholami et al. 2022), pruning (LeCun, Denker, and Solla 1989; Frankle and Carbin 2018; Liu et al. 2018; Cheng, Zhang, and Shi 2024), knowledge distillation (Hinton 2015; Heo et al. 2019; Wang and Yoon 2021), and neural architecture search (Zoph 2016; Elsken, Metzen, and Hutter 2019), as surveyed in (Cheng et al. 2018; Deng et al. 2020). Among these, quantization has emerged as the most actively studied technique. It serves as an effective compression method by reducing unnecessary representational ranges in the model. However, it often requires fine-tuning or calibration to match full-precision (FP) model performance (Gholami et al. 2022). To address this, zero-shot quantization (ZSQ), also known as data-free quantization, has been proposed to compress models without the original training data.
@@
ZeroQ (CVPR 20)
GDFQ (ECCV 20)
DSG (CVPR 21)
MixMix (CVPR 21)
Qimera (NeurIPS 21)
IntraQ (CVPR 22)
GENIE (CVPR 23)
AdaDFQ (CVPR 23)
Casual-DFQ (ICCV 23)
TexQ (NeurIPS 23)
RIS (AAAI 24)
GenQ (ECCV 24)
SQuant (ICLR 22)
UDFC (ICCV 23)
AIT (CVPR 22)
PLF (CVPR 24)
SynQ (ICLR 25)
@@
Table 1: Categorizing ZSQ algorithms. PTQ denotes post-training quantization, and QAT denotes quantization-aware training. Data generation methods typically support both PTQ and QAT, whereas violet-marked methods utilize fine-tuning for data generation and support only QAT.
Since the introduction of ZeroQ (Cai et al. 2020), studies on ZSQ have advanced in two main directions. The first direction focuses on data generation, where synthetic data are created from the FP model. The second direction focuses on effectively applying the activation distributions of the synthetic data to the quantized (Q) model. This second direction is further divided into post-training quantization (PTQ), which calibrates the activation distributions, and quantization-aware training (QAT), which fine-tunes the Q model directly. We categorize existing ZSQ methods based on these directions, as summarized in Table 1.
(a) Layer-wise Quantization
(b) GranQ (Channel-wise, Ours)
First, data generation studies focus on generating high-quality data to effectively train the Q model (Xu et al. 2020; Qian et al. 2023; Li et al. 2024). Meanwhile, calibration (PTQ) studies aim to minimize the quantization error by calibrating the Q model without additional training (Guo et al. 2022; Bai et al. 2023). Finally, fine-tuning (QAT) studies focus on transferring key information from the FP model to the Q model through knowledge distillation (Choi et al. 2022; Fan et al. 2024; Hong et al. 2025; Kim, Kim, and Kang 2025).
Based on this analysis, we introduce GranQ, a novel ZSQ method that achieves efficient per-channel quantization via vectorized pre-scaling of input-dependent activations. This dynamic adjustment minimizes activation loss and preserves the original activation values by reducing quantization errors, as shown in Figure 1. The proposed method effectively handles activation loss in low-bit quantization and achieves state-of-the-art (SOTA) performance in QAT settings on the CIFAR and ImageNet datasets. Furthermore, we apply vectorization to the pre-scaling step, which is typically omitted in conventional channel-wise quantization, thereby reducing latency and enabling fine-grained activation quantization. Our contributions can be summarized as follows:
  • We propose GranQ, a novel method that supports granular quantization and maintains computational efficiency. Although per-channel activation quantization is known to improve precision, it has not been widely adopted due to its high computational cost. We address this by introducing vectorized pre-scaling, which integrates per-channel scaling into the quantization step, allowing accumulation to proceed efficiently without runtime scaling overhead. To the best of our knowledge, our approach is the first to address the ZSQ problem.
Related Work
Quantization
Quantization reduces the representational range of deep neural networks (DNNs) and minimizes memory usage. It is typically categorized into post-training quantization (PTQ) and quantization-aware training (QAT) (Gholami et al. 2022; Banner, Nahshan, and Soudry 2019). PTQ applies quantization after training without further updates, which makes it efficient and easy to implement. However, it is sensitive to scaling errors and often relies on calibration to improve accuracy. In contrast, QAT incorporates quantization during training and optimizes quantized activations using the straight-through estimator (Gholami et al. 2022; Yin et al. 2019). Both methods require data. PTQ uses it for calibration, while QAT uses it for fine-tuning.. In practice, access to training data is often limited. To address this, zero-shot quantization (ZSQ) has been proposed to perform quantization without using original data (Xu et al. 2020).
Zero-shot Quantization
@@
As summarized in Table 1, ZSQ was initially introduced by ZeroQ (Cai et al. 2020), leading to the development of various ZSQ algorithms. These studies have explored diverse approaches to improve quantization performance under data-free settings. Recent studies have focused on generating high-quality synthetic data (Jeon, Lee, and Kim 2023; Li et al. 2024) and developing more effective methods to utilize them (Choi et al. 2022; Hong et al. 2025; Kim, Kim, and Kang 2025). However, current ZSQ methods heavily rely on augmented synthetic inputs. This causes large variation in activation scales across channels. In practice, such input-dependent channel-wise quantization is rarely adopted due to its high computational cost. Therefore, recent studies have applied channel-wise activation quantization either in a limited conditions within QAT settings (Cho and Yoo 2020), or exclusively in PTQ scenarios with access to a small amount of data (Yvinec et al. 2023; Xu et al. 2024). This highlights the importance of efficient channel-wise activation quantization in QAT.
Preliminaries and Problem Definition
@@
Activation Quantization
Activation quantization reduces the precision of intermediate activation values by converting them into low-bitwidth integers (Choi et al. 2018). It is commonly used with weight quantization to compress models, and the linear quantization scheme is most widely adopted. This process typically consists of quantization, scaling parameter computation, and dequantization (Hubara et al. 2018; Gholami et al. 2022; Fang et al. 2020).
@@
where $x_{\min}$ and $x_{\max}$ denote the range of activation values, and $b$ is the bit-width. A large range yields a large $s$, increasing quantization error, a smaller range allows finer resolution. As shown below, in quantized models, per-channel scaling is embedded directly into the accumulation path, requiring each channel to be individually scaled during computation:
$y_{l}=\sum_{c=1}^{C}w_{c}\cdot s_{c}\cdot(x_{q,c}-z_{c})$ (3)
However, applying the per-channel scaling factor $s_{c}$ inside the accumulation loop introduces computational overhead, making efficient integer-domain execution difficult. This scaling bottleneck hinders parallelism in core operations such as convolution and matrix multiplication, reducing the efficiency of quantized models (Nagel et al. 2021).
Problem Definition
@@
(Problem 2) Accumulation Bottleneck in Channel-wise Quantization The main challenge in channel-wise quantization lies not in quantizing activations but in the repeated per-channel scaling during accumulation. This disrupts vectorized execution and significantly increases computational overhead.
Observation and Methodology
We observe the causes of activation distortion during quantization and analyze how to mitigate them. Based on this, we introduce our method, GranQ.
Observation
@@
Table 2: Average cosine similarity and relative error for 3-bit activation quantization on ResNet-20 (CIFAR-100). GranQ shows lower activation distortion than layer-wise, with higher similarity and lower error.
[…]
Furthermore, Figure 1(a) illustrates how layer-wise quantization leads to coarse activation representations across layers. Fine-grained activation values are often mapped to limited discrete levels, resulting in a loss of detail and reduced representational capacity. This degradation contributes to decreased inference performance.
Methodology
Based on our analysis, we propose GranQ, a fine-grained ZSQ method designed to reduce computational overhead, as illustrated in Figure 2. GranQ applies channel-wise scaling through activation decomposition, and achieves fast parallel processing via pre-scaling for parallel accumulation, which enables efficient computation of both scaling and quantization.
Algorithm 1 GranQ: Pre-scaling for Parallel Accumulation
1:Activation tensors $\{A_{l}\}_{l=1}^{L}$, where $A_{l}\in\mathbb{R}^{C\times H\times W}$, quantization bit $b$, weight tensors $\{w_{l,c}\}_{l=1,c=1}^{L,C}$
2:Output vectors $\{y_{l}\}_{l=1}^{L}$
3:for each layer $l$ in $L$ do
4:  $A_{l}\leftarrow\text{reshape}(A_{l})\in\mathbb{R}^{C\times(HW)}$
5:  $\vec{A}_{\min},\vec{A}_{\max}\leftarrow\min_{h,w}(A_{l}),\max_{h,w}(A_{l})$
6:  $\vec{s}_{l}\leftarrow(\vec{A}_{\max}-\vec{A}_{\min})/(2^{b}-1)$
7:  $\vec{z}_{l}\leftarrow\left\lfloor-\vec{A}_{\min}/\vec{s}_{l}\right\rceil$
8:  $\vec{A}_{q,l}\leftarrow\left\lfloor A_{l}/\vec{s}_{l}+\vec{z}_{l}\right\rceil$ $\triangleright$ quantize (int mapping)
9:  $\vec{\tilde{A}}_{l,c}\leftarrow\vec{s}_{l}\cdot(\vec{A}_{q,l}-\vec{z}_{l})$ $\triangleright$ dequantize & pre-scale
10:  Compute output:
$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\vec{\tilde{A}}_{l,c}$
11:end for
12:return $\{y_{l}\}_{l=1}^{L}$
Activation Decomposition
We propose activation decomposition, which reshapes each activation map from a three-dimensional tensor of shape $(C\times H\times W)$ into a two-dimensional matrix of shape $(C\times HW)$. This transformation enables decomposition of the activation map along the channel axis, allowing each channel to be processed independently. It also facilitates vectorized computation of per-channel statistics (e.g., min and max) required for scaling factor calculation during quantization. We define the activation decomposition as shown in Equation 4.
@@
Pre-scaling for Parallel Accumulation
@@
$y_{l}=\sum_{c=1}^{C}w_{l,c}\cdot\vec{\tilde{A_{l,c}}},\quad\text{where }\vec{\tilde{A_{l,c}}}=\vec{s}_{c}\cdot\left(\left\lfloor\frac{\vec{A}_{l,c}}{\vec{s}_{c}}+\vec{z}_{c}\right\rceil-\vec{z}_{c}\right)$ (6)
In the pre-scaling for parallel accumulation stage, the scaling factors $\vec{s}_{c}$ and zero-points $\vec{z}_{c}$ are computed in a vectorized form across channels to dynamically adapt to the diverse activation distributions induced by synthetic inputs. This contrasts with conventional per-channel quantization, which applies fixed scaling parameters per channel regardless of input variation.
As shown in Equation 6, each activation $\vec{A}_{l,c}$ is first quantized using its corresponding vectorized scaling and zero-point, and then dequantized to reconstruct the floating-point value. The resulting $\vec{\tilde{A_{l,c}}}$ represents a pre-scaled activation that is already adjusted for accumulation. By applying vectorized scaling, quantization, and dequantization uniformly across channels, this process enables efficient integer-domain computation and parallel accumulation with weights $w_{l,c}$.
Experiments
@@
Experimental Setup and Details
The experiments were conducted using widely adopted ZSQ evaluation datasets, including CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), and ImageNet (ILSVRC 2012) (Deng et al. 2009) validation datasets. For the CIFAR datasets, ResNet-20 (He et al. 2016) was used as the quantization model, whereas ResNet-18 (He et al. 2016), ResNet-50 (He et al. 2016), and MobileNetV2 (Sandler et al. 2018) were employed for ImageNet. All experiments were conducted using the SGD optimizer (Ruder 2016) with a momentum of 0.9 and weight decay of 1e-4. The CIFAR-10 and CIFAR-100 experiments were each conducted for 200 epochs, with batch sizes of 16 and 200, respectively. For ImageNet, we trained for 400 epochs with a batch size of 16. The initial learning rate was set to 1e-4 for CIFAR-10 and CIFAR-100, and 1e-5 for ImageNet, with multi-step learning rate decay applied. The decay steps were set to 100, 200, and 300 epochs for CIFAR, and at 350 and 400 epochs for ImageNet, with a decay rate of 0.1. We compared our method with existing ZSQ methods (Xu et al. 2020; Choi et al. 2022; Qian et al. 2023; Chen et al. 2024; Bai et al. 2024; Fan et al. 2024; Li et al. 2024; Hong et al. 2025; Kim, Kim, and Kang 2025). For data generation, we followed the AdaDFQ (Qian et al. 2023) approach based on ACGAN (Odena, Olah, and Shlens 2017). Further implementation details are provided in the appendix, Section A and A. In our ablation study, layer-wise quantization was applied to all layers containing activation functions, while channel-wise quantization was performed per channel at the batch level.
@@
Cifar-10
ResNet-20
3ww3aa 75.11 - 84.89 86.47 - - 86.76 88.11 91.37
4ww4aa 90.25 90.49 92.31 92.68 92.59 - 92.64 92.76 93.52
5ww5aa 93.38 92.98 93.81 - 93.59 - 93.83 - 94.06
Cifar-100
ResNet-20
3ww3aa 47.61 - 52.74 55.87 - - 54.68 57.28 62.73
4ww4aa 63.39 61.05 66.81 67.18 65.99 - 66.94 67.34 68.79
5ww5aa 66.12 68.40 69.93 - 69.55 - 69.75 - 70.05
ImageNet
ResNet-18
3ww3aa 20.23 - 38.10 50.28 - 68.18 49.88 52.02 64.41
4ww4aa 60.60 65.73 66.53 67.73 67.55 70.03 65.89 67.90 70.39
5ww5aa 68.49 70.28 70.29 - 70.59 - 69.40 - 71.31
MobileNetV2
3ww3aa 1.46 - 28.99 32.80 - 59.15 30.56 34.21 62.42
4ww4aa 59.43 66.47 65.41 67.07 - 69.65 64.85 67.27 70.62
5ww5aa 68.11 71.96 71.61 - - - 71.71 - 72.49
ResNet-50
3ww3aa 0.31 - 17.63 25.27 - 73.99 24.50 26.89 70.76
4ww4aa 54.16 68.27 68.38 70.72 71.54 76.10 68.75 71.05 76.63
5ww5aa 71.63 76.00 76.03 - 76.36 - 75.90 - 77.58
@@
CIFAR-10/100. As indicated in Table 3, GranQ consistently achieved the highest accuracy across all bitwidths. For CIFAR-10, it attained accuracies of 94.06% (5-bit), 93.52% (4-bit), and 91.37% (3-bit). For CIFAR-100, the results were 70.05% (5-bit), 68.79% (4-bit), and 62.73% (3-bit). Notably, GranQ outperformed SynQ (Kim, Kim, and Kang 2025), the previous SOTA, by +5.45% in the CIFAR-100 3-bit setting. This result demonstrates its ability to effectively overcome the limitations of conventional low-bitwidth quantization techniques. Overall, GranQ consistently outperformed existing methods across all bitwidths in both CIFAR-10 and CIFAR-100, with particularly strong improvements in the 3-bit quantization setting. Remarkably, in the CIFAR-10 5-bit setting, GranQ even surpassed the performance of the FP model. These results suggest that we can effectively apply GranQ to small- and medium-scale datasets.
ImageNet. In the ImageNet experiments, GranQ achieved competitive performance across various bitwidths. Specifically, for ResNet-18, it attained top accuracies of 70.39% (4-bit) and 71.31% (5-bit). For ResNet-50, it achieved the highest accuracies of 76.63% (4-bit) and 71.31% (5-bit). Additionally, in the MobileNetV2 setting, GranQ achieved SOTA performance across all bitwidths (3, 4, and 5-bit). In the 3-bit setting, GranQ achieved the second-best accuracies, with 64.41% on ResNet-18 and 70.76% on ResNet-50. These results are -3.77% and -3.23% lower than those of GenQ, respectively. However, GenQ uses a pre-trained diffusion model (Ho, Jain, and Abbeel 2020) and prompts during data generation, leading to slower speeds compared with those of typical GAN-based methods (Xu et al. 2020). In contrast, GranQ enhances the quantization mechanism without heavily relying on diffusion-based data generation, as GenQ does. Although GenQ demonstrates strong performance, its code is not publicly available, and thus we could not include it in our experiments. Nevertheless, since our method is compatible with existing QAT frameworks, it could potentially benefit from GenQ’s data generation if integration becomes possible.
Ablation Study
Effectiveness Evaluation
@@
First, we analyzed the impact of GranQ on data synthesis-based quantization methods, specifically GDFQ (Xu et al. 2020) and AdaDFQ (Qian et al. 2023). The integration of GranQ into GDFQ (Xu et al. 2020) led to a notable improvement, with accuracy rising from 47.61% to 59.04% (+11.43%) in the 3-bit setting and from 63.39% to 66.97% (+3.58%) in the 4-bit setting. Similarly, AdaDFQ (Qian et al. 2023) exhibited improvements of +9.99% in the 3-bit setting (from 52.74% to 62.73%) and +1.98% in the 4-bit setting (from 66.81% to 68.79%). These results suggest that GranQ effectively reduces quantization errors when combined with data synthesis-based quantization methods, leading to enhanced model performance.
Furthermore, GranQ also exhibits significant improvements in methods focused on training quantized models, such as Qimera+AIT (Choi et al. 2022) and AdaDFQ+AKT (Hong et al. 2025). For Qimera+AIT (Choi et al. 2022), the 3-bit accuracy increased from 45.70% to 60.42% (+14.72%). Similarly, for AdaDFQ+AKT (Hong et al. 2025), the accuracy improved from 54.68% to 62.01% (+7.33%). These findings demonstrate that GranQ is not only effective in data synthesis-based methods but also enhances performance during the model training process.
Efficiency Evaluation
@@
Training
Time (sec/epoch)
Layer-wise 49.98 0.227 10.92
Channel-wise (Baseline) 62.68 103.671 17370.54
Channel-wise (GranQ) 62.73 0.236 12.43